Given an RC series circuit that has an emf source of 50 volts, a resistance of 20k ohms, a capacitance of 6 microfarad and the initial charge of the capacitor is 1 microcoulomb. What is the charge in the capacitor at the end of 0.01 second? What is the current in the circuit at the end of 0.05 seconds?


1
Expert's answer
2021-11-29T19:29:16-0500
"Rq'+\dfrac{1}{C}q=V"

"2\cdot10^{4}q'+\dfrac{1}{6\cdot10^{-6}}q=50"

"q'+\dfrac{25}{3}q=0.0025"

"q'=-\dfrac{25}{3}(q-0.0006)"

"\dfrac{dq}{q-0.0003}=-\dfrac{25}{3}dt"

Integrate


"q=0.0003+c_1e^{-25t\/3}"

Given "q(0)=10^{-6}"

"0.000001=0.0003+c_1"

"c_1=-0.000299"

"q(t)=0.0003-0.000299e^{-25t\/3}"

"q(0.01)=0.0003-0.000299e^{-25(0.01)\/3}"

"=24.91\cdot10^{-6} (C)"

"24.91\ \mu C"

"i(t)=q'(t)=-0.000299(-\dfrac{25}{3})e^{-25t/3}"

"=\dfrac{0.007475}{3}e^{-25t/3}"

"i(0.05)=\dfrac{0.007475}{3}e^{-25(0.05)/3}=0.0016426(A)"

"=1.64(\ mA)"

"1.64\ mA"